OpenStudy (anonymous):

I think my question got lost - is there an easy was of finding the largest perfect square factor?

OpenStudy (anonymous):

\[\sqrt{\frac{ 148 }{ 144? }}\]

OpenStudy (anonymous):

Yes, It depends on the situation

OpenStudy (anonymous):

what situation and how

OpenStudy (anonymous):

The answer would become \[\sqrt{37}\]

OpenStudy (anonymous):

Divided by 6

OpenStudy (anonymous):

square root 37/6

OpenStudy (anonymous):

? how did you get the 37

OpenStudy (anonymous):

Example: what is √20 ? Children first learn to find the easy square roots that are whole numbers, but quickly the question arises as to what are the square roots of all these other numbers. You can start out by noting that (dealing here only with the positive roots) since √16 = 4 and √25 = 5, then √20 should be between 4 and 5 somewhere. Then is the time to make a guess, for example 4.5. Square that, and see if the result is over or under 20, and improve your guess based on that. Repeat the process until you have the desired accuracy (amount of decimals). It's that simple and can be a nice experiment for children.

OpenStudy (anonymous):

thanks but I still don't get it

OpenStudy (anonymous):

A square root is the most common root operation. A root is the inverse operation of an exponent (which means that it undoes an exponential operation), and so a square root is an operation that undoes an exponent of 2. For example, You can read the symbol either as “the square root of” or as “radical.” So, read as either “the square root of 9” or “radical 9.” As you can see, when you take the square root of any square number, the result is the number that you multiplied by itself to get that square number in the first place. For example, to find you ask the question, “What number when multiplied by itself equals 100?” The answer in this case is 10, because